### v2

# Environmental Breakdown of Topological Interface States in Armchair Graphene Nanoribbon Heterostructures

Abstract
We theoretically investigate the stability and transport properties of topological interface states (IFs) in 9-7-9 and 15-13-15 armchair graphene nanoribbon heterostructures (AGNRHs) laterally embedded in boron nitride (BN) sheets. Two configurations, $n$-BNNR/AGNRH/$n$-BNNR and $n$-BNNR/AGNRH/$n$-NBNR, corresponding to same-topology and reverse-topology BN environments, are examined within a tight-binding framework. Using a bulk boundary perturbation approach, we show that in BNNR/AGNRH/BNNR the IFs are destroyed by chirality breaking induced by symmetric BN environments at both interfaces. In contrast, the IFs in the reverse-topology structure remain robust against lateral interface interactions from BN atoms. Transport calculations further demonstrate that the surviving IFs in BNNR/AGNRH/NBNR exhibit the characteristic behavior of topological double quantum dots, with an enhanced interdot hopping strength compared with vacuum boundary conditions. These results reveal that BN environments can either suppress or reinforce topological interface states, depending critically on the topology of the surrounding nanoribbons.
I Introduction
Bottom-up synthesis techniques enable the fabrication of graphene nanoribbons (GNRs) with atomic precision [Cai, 1DJRizzo, 10]. Compared with zigzag GNRs (ZGNRs), armchair GNRs (AGNRs) exhibit tunable electronic phases [LlinasJP, 5], making them promising candidates for nanoelectronic applications [WangHM, 9]. More complex GNR architectures, including 9-7-9 and 7-9-7 armchair graphene nanoribbon heterostructures (AGNRHs) [DJRizzo, 10] and Janus GNR segments [SongST, 11], have also been experimentally realized. Scanning tunneling microscopy measurements have confirmed the presence of topological states (TSs) in these systems, such as end states (ESs) and interface states (IFs) in 9-7-9 and 7-9-7 AGNRHs [DJRizzo, 10].
Although topological IFs in AGNRHs, such as 9-7-9 ribbons, have been extensively studied in idealized or substrate-agnostic models [DJRizzo, 10,DavidKuo, 12,DavidKuo1, 13], the role of the environment particularly lateral embedding materials that break sublattice symmetry remains largely unexplored. In realistic device configurations, AGNRHs are inevitably coupled to surrounding materials, raising a critical question: are topological interface states robust against such environmental perturbations ? Here, we theoretically investigate a 9-7-9 AGNRH laterally embedded in boron nitride nanoribbons (BNNRs), as illustrated in Fig. 1(a) and 1(b). Two distinct configurations are considered, in which the upper and lower BNNRs realize either the same-topology or reverse-topology scenario. Owing to the difficulty of employing density functional theory (DFT) to compute transmission coefficients in large 9-7-9-AGNRH/BNNR junctions [SonYW, 14NikitaVT, 16], we adopt a tight-binding model combined with the Green function technique to evaluate quantum transport in AGNRHs embedded in BN environments [QiaoZH, 17,JungJ, 18].
Surprisingly, we find that BNNR/AGNRH/NBNR structures with reverse topology preserve the functionality of topological double quantum dots (TDQDs), in which a single energy level associated with each quantum dot remains well isolated from the bulk states. Moreover, the electron hopping strength between the TDQDs is significantly enhanced compared with the vacuum boundary condition. This behavior is highly unusual when contrasted with conventional semiconductor DQD systems [Ono, 19ZajacDM, 26], whose operation is typically restricted to ultra-low temperatures due to the small energy separation between bound and continuum states. Nanoelectronic devices based on TDQDs therefore offer a promising route toward robust high-temperature quantum functionality.
Figure 1: Schematic illustration of an armchair graphene nanoribbon heterostructure (AGNRH) laterally embedded in a hexagonal boron nitride (BN) sheet, forming two distinct configurations: (a) a same-topology structure, 2-BNNR/9-7-9-AGNRH/2-BNNR, and (b) a reverse-topology structure, 2-BNNR/9-7-9-AGNRH/2-NBNR. White and blue sites denote boron and nitrogen atoms, respectively. $R_{1}$ and $R_{2}$ represent the unit cells (u.c.) of the 2-BNNR/9-AGNR/2-BNNR and 3-BNNR/7-AGNR/3-BNNR junctions, respectively. The tunable parameter $t_{in}$ characterizes the interfacial coupling between the AGNRH and the BNNRs. The symbols $\Gamma_{L}$ ($\Gamma_{R}$) represent the electron tunneling rates between the left (right) electrode and the leftmost (rightmost) atoms at the zigzag edges.
II Calculation Methodology
To investigate the transport properties of the BNNR/AGNRH/NBNR heterojunction connected to metallic electrodes, we employ a combination of a tight-binding model and the Green’s function technique [QiaoZH, 17]. The system Hamiltonian, illustrated in Fig. 1(b), is written as $H=H_{0}+H_{GNR}$, where $H_{0}$ describes the electrodes and their coupling to the nanoribbon system, and $H_{GNR}$ represents the BNNR/AGNRH/NBNR structure.
$H_{0}$ $=$ $\sum_{k}\epsilon_{k}a^{\dagger}_{k}a_{k}+\sum_{k}\epsilon_{k}b^{\dagger}_{k}b_{k}$
$+$ $\sum_{\ell}\sum_{k}V^{L}_{k,\ell,j}d^{\dagger}_{\ell,j}a_{k}+\sum_{\ell}\sum_{k}V^{R}_{k,\ell,j}d^{\dagger}_{\ell,j}b_{k}+h.c.$
The first two terms describe free electrons in the left and right metallic electrodes. The operators $a^{\dagger}_{k}$ ($b^{\dagger}_{k}$) create an electron with momentum $k$ and energy $\epsilon_{k}$ in the left (right) electrode, respectively. The coupling amplitudes $V^{L}_{k,\ell,j=1}$ and $V^{R}_{k,\ell,j=M}$ describe tunneling between the left (right) electrode and its adjacent atom located in the $\ell$-th row at the zigzag edge.
$H_{GNR}$ $=$ $\sum_{\ell,j}E_{\ell,j}d^{\dagger}_{\ell,j}d_{\ell,j}$
- $\sum_{\ell,j}\sum_{\ell^{\prime},j^{\prime}}t_{(\ell,j),(\ell^{\prime},j^{\prime})}d^{\dagger}_{\ell,j}d_{\ell^{\prime},j^{\prime}}+h.c,$
Here, $E_{\ell,j}$ denotes the on-site energy of the orbital at site $(\ell,j)$, while $d^{\dagger}_{\ell,j}$ and $d_{\ell,j}$ are the electron creation and annihilation operators, respectively. The parameter $t_{(\ell,j),(\ell^{\prime},j^{\prime})}$ characterizes the hopping amplitude between sites $(\ell^{\prime},j^{\prime})$ and $(\ell,j)$. For the BNNR/AGNRH/NBNR system, we assign on-site energies $E_{B}=\Delta_{B}$, $E_{N}=-\Delta_{N}$, and $E_{C}=0$ for boron, nitrogen, and carbon atoms, respectively. Variations in hopping strengths between different atomic species are neglected, and the condition $E_{B}=-E_{N}=\Delta=2.7$ eV is adopted for simplicity [DingY, 27KuoJL, 30]. The nearest-neighbor hopping energy is set to $t_{(\ell,j),(\ell^{\prime},j^{\prime})}=t_{pp\pi}=2.7$ eV.
To quantify interfacial interactions induced by the BN environment, we introduce an inter-nanoribbon hopping parameter $t_{in}$ based on bulk boundary perturbation approach[DavidKuo, 12], as illustrated in Fig. 1(a). When $t_{in}=0$, the BNNR/AGNRH/BNNR (or BNNR/AGNRH/NBNR) structure is decoupled into three isolated segments–the lower BNNR, the central 9-7-9-AGNRH, and the upper BNNR (or NBNR)–each subject to vacuum boundary conditions. Varying $t_{in}$ effectively models different strengths of lateral interface interactions arising from surrounding two-dimensional insulating environments [PerepichkaDF, 31].
The bias-dependent transmission coefficient ${\cal T}_{GNR}(\varepsilon)$ is calculated using ${\cal T}_{GNR}(\varepsilon)=4Tr[\Gamma_{L}(\varepsilon)G^{r}(\varepsilon)\Gamma_{R}(\varepsilon)G^{a}(\varepsilon)]$ [DavidKuo1, 13], where $\Gamma_{L}(\varepsilon)$ and $\Gamma_{R}(\varepsilon)$ denote the tunneling rate (in energy units) at the left and right leads, respectively, and $G^{r}(\varepsilon)$ and $G^{a}(\varepsilon)$ are the retarded and advanced Green’s functions of the GNRs, respectively. In terms of tight-binding orbitals, $\Gamma_{\alpha}(\varepsilon)$ and Green’s functions are matrices. The expression for $\Gamma_{L(R)}(\varepsilon)$ is derived from the imaginary part of the self-energies, denoted as $\Sigma^{r}_{L(R)}(\varepsilon)$, and is given by $\Gamma_{L(R)}(\varepsilon)=-\text{Im}(\Sigma^{r}_{L(R)}(\varepsilon))=\pi\sum_{k}|V^{L(R)}_{k,\ell,j=1(M)}|^{2}\delta(\varepsilon-\epsilon_{k})$. In the wide-band limit, $\Gamma_{L(R)}(\varepsilon)$ is replaced by an energy-independent constant matrix $\Gamma_{L(R)}$ [DavidKuo, 12].
III Results and discussion
Interface-induced modifications of AGNRs embedded in BN sheets have been experimentally observed and shown to alter the band gaps between conduction and valence subbands [Levendorf, 32GengDC, 35]. To elucidate how BN environments affect the topological IFs of 9-7-9 -AGNRHs, we first examine the influence of BNNRs with different topologies on individual 7-AGNR and 9-AGNR with $R_{1}$ unit cells. Figures 2(a) and 2(b) show the calculated band structures of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-AGNR/4-BNNR, corresponding to same-topology configurations. The resulting band gaps are $1.22$ eV and $1.19$ eV, respectively, compared with $1.0$ eV and $0.9$ eV for isolated 7-AGNR and 9-AGNR under vacuum boundary conditions.
For same-topology BN lateral embedding, both the lower and upper BN edges induce sublattice potentials of the same sign. Consequently, the boundary condition imposed by BN resembles that of a vacuum (infinite potential barrier), except that the BN case corresponds to a finite potential barrier. This explains why the band-gap differences between vacuum and BN boundary conditions are relatively small. Indeed, when $\Delta$ is set to a sufficiently large value (e.g., $\Delta=100~\mathrm{eV}$), the two results become nearly indistinguishable. These findings indicate that BN lateral embedding introduces only weak perturbations to the bulk electronic structure when the surrounding BNNRs share the same topology [JungJ, 18,KuoJL, 30].
In contrast, for inverse-topology configurations shown in Figs. 2(c) and 2(d), the band gaps of 4-BNNR/7-AGNR/4-NBNR and 4-BNNR/9-AGNR/4-NBNR are significantly reduced to $0.394$ eV and $0.454$ eV, respectively. This pronounced gap renormalization highlights the strong influence of the BN environment when the lower BNNR has an opposite topology to the upper NBNR. In the inverse-topology case, the sublattice potentials induced at the lower and upper BN edges have opposite signs. Such a boundary condition differs fundamentally from that of AGNRs under vacuum boundary conditions, leading to substantial modifications of the band-edge states.
These tight-binding results are in good agreement with previous DFT calculations [JungJ, 18,KuoJL, 30]. The small residual discrepancies suggest that higher-order hopping processes play a negligible role in determining the low-energy conduction and valence subbands. Furthermore, we find that varying the BNNR width ($n=6,8,10,\ldots$) does not qualitatively modify the band-edge dispersions, thereby justifying the use of 4-BNNRs in the following analysis.
Figure 2: Quasi-1D electronic band structures of AGNRs embedded in hexagonal BN sheets. Same-topology configurations: (a) 4-BNNR/7-AGNR/4-BNNR and (b) 4-BNNR/9-AGNR/4-BNNR. Inverse-topology configurations: (c) 4-BNNR/7-AGNR/4-NBNR and (d) 4-BNNR/9-AGNR/4-NBNR. We have adopted $t_{in}=2.7$ eV and $\Delta=2.7$ eV. $L$ is the length of $R_{1}$ unit cell.
In realistic experiments, AGNR and AGNRH segments possess finite lengths, typically below $20$ nm[DJRizzo, 10]. We therefore examine the discrete energy spectra of finite 7-AGNR and 9-7-9 AGNRH segments embedded in BN environments with same and inverse topology, as shown in Fig. 3. Figures 3(a) and 3(b) display the energy levels of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-7-9-AGNRH/4-BNNR as functions of the interfacial hopping strength $t_{in}$ for the same-topology case. Due to the finite size of the 4-BNNR/7-AGNR/4-BNNR segment $(N=15,M=80)$, two in-gap energy levels, denoted as $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$, emerge from the ESs, as shown in Fig. 3(a). In the absence of BN coupling ($t_{in}=0$), these ESs form near-zero-energy modes. As $t_{in}$ increases, $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$ shift markedly away from zero energy, whereas the bulk-derived states remain largely unchanged. This behavior is consistent with the infinite-length band structures shown in Fig. 2, indicating that BN coupling primarily affects boundary-localized states in the same-topology case.
Figure 3(b) shows four in-gap energy levels originating from two ESs and two IFs in the 9-7-9 AGNRH segment. The IFs arise from the effective end states of the central 7-AGNR segment with $R_{2}$ unit cells, as discussed in our previous work [DavidKuo, 12]. Within the physically relevant range $0.8~t\leq t_{in}\leq t$, corresponding to realistic BN coupling strengths, the IF levels $\Sigma_{IF,C(V)}$ are no longer well isolated from the bulk states. This behavior contrasts sharply with the $t_{in}=0$ limit and reflects the enhanced sublattice polarization induced by the symmetric BN environment, which drives the IF levels toward the bulk continuum.
The situation changes dramatically in the inverse-topology configuration shown in Figs. 3(c) and 3(d), where finite 7-AGNR and $9_{6}$-$7_{8}$-$9_{6}$ AGNRH segments are embedded between a BNNR and an NBNR. Although the bulk band gap ($E_{g}=E_{C}-E_{V}$) is substantially reduced for $0.8~t\leq t_{in}\leq t$, the energy levels associated with the ESs and IFs in the 9-7-9 AGNRH remain nearly unchanged and well separated from the bulk continuum. Such robustness is highly nontrivial and indicates that the inverse-topology configuration effectively preserves the IFs against strong environmental perturbations.
Figure 3: Energy spectra of finite AGNR (AGNRH) segments as functions of interfacial hopping $t_{in}$ at $\Delta=2.7$ eV. Same-topology scenario: (a) 4-BNNR/7-AGNR/4-BNNR and (b) 4-BNNR/9-7-9-AGNRH/4-BNNR. Inverse-topology scenario: (c) 4-BNNR/7-AGNR/4-NBNR and (d) 4-BNNR/9-7-9-AGNRH/4-NBNR.
To clarify the contrasting behaviors observed in Figs. 3(b) and 3(d), we plot the charge densities of $\Sigma_{IF,C}$ for different $t_{in}$ values in Fig. 4. Figures 4(a)-4(c) and Figs. 4(d)-4(f) correspond to the inverse-topology and same-topology configurations, respectively. For $t_{in}=0$, shown in Fig. 4(a), the AGNRH is completely decoupled from the BN environment, corresponding to vacuum boundary conditions. The state $\Sigma_{IF,C}$ originates from the coupling between a 9-7 junction IF state with A-sublattice chirality ($\Psi_{IF,A}(t_{in}=0)$) and a 7-9 junction IF state with B-sublattice chirality ($\Psi_{IF,B}(t_{in}=0)$). When a weak coupling $t_{in}=0.2~t$ is introduced, the charge density of $\Sigma_{IF,C}$ partially penetrates into the BN regions [Fig. 4(b)] while preserving mirror symmetry across the ribbon. Even for strong coupling $t_{in}=0.8~t$, as shown in Fig. 4(c), $\Sigma_{IF,C}$ maintains nearly equal contributions from $\psi_{IF,A}(t_{in}\neq 0)$ and $\psi_{IF,B}(t_{in}\neq 0)$. The charge-density distribution remains mirror-inverted across the ribbon due to the global inversion symmetry of the system. The increased energy separation between $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ is mainly attributed to the enhanced overlap between the wave functions $\psi_{IF,A}$ and $\psi_{IF,B}$. For finite $t_{in}$, the decay lengths of $\psi_{IF,A}$ and $\psi_{IF,B}$ increase as a result of interfacial-induced delocalization. Notably, charge densities simultaneously appear on both A- and B-sublattice sites at rows $i=5$ and $i=13$, reflecting the symmetric hybridization between the two junction states.
In contrast, Figs. 4(d)-4(f) demonstrate that in the same-topology configuration the IF charge densities become strongly polarized even at $t_{in}=0.2~t$, with dominant contributions from $\psi_{IF,A}$. The A(B)-sublattice IF states couple to nitride (boron) atoms at the lower and upper BN edges, respectively, leading to the formation of new bonding and antibonding states [PrunedaJM, 28]. As a result, $\Sigma_{IF,C(V)}$ shifts above (below) the mid-gap energy. This behavior indicates that the symmetric BN environment generates an effective field along the armchair direction, thereby breaking the inversion symmetry between the 9-7 and 7-9 junctions.
To further elucidate the influence of BN sheets on the IFs shown in Figs. 3 and 4, we introduce an effective Hamiltonian for the IF subspace:
$H_{eff,IF}$ (3)
$=$ $\left[\begin{array}[]{cc}E_{L}+\Sigma_{A}&t_{eff,LR}\\ t_{eff,LR}&E_{R}+\Sigma_{B}\\ \end{array}\right],$ (6)
where $E_{L}=0$ and $E_{R}=0$ denote the energy levels of the 9-7 and 7-9 junction IF states in the absence of $t_{in}$. $\Sigma_{A}$ and $\Sigma_{B}$ represent the self-energies induced by the interfacial hopping $t_{in}$.
For the inverse-topology configuration, the effective Hamiltonian reduces to
$H_{eff,IF,I}$ (7)
$=$ $\left[\begin{array}[]{cc}0&t_{eff,LR}\\ t_{eff,LR}&0\\ \end{array}\right],$ (10)
since the self-energies satisfy $\Sigma_{A(B)}=0$. More specifically, $\Sigma_{A(B)}=t_{in}^{2}(G_{Bo}+G_{Ni})$, where $G_{Bo}$ and $G_{Ni}$ are the Green’s functions describing electron propagation on boron and nitride atoms, respectively. Under the approximation $\Delta_{B}=-\Delta_{N}$, one has $G_{Bo}\approx-G_{Ni}$, leading to nearly vanishing self-energies. In the inverse-topology configuration, the upper BN ribbon is flipped relative to the lower one, reversing the BN sublattice alignment at the two interfaces and restoring chiral symmetry in the IF subspace.
For the same-topology configuration, the effective Hamiltonian becomes
$H_{eff,IF,S}$ (11)
$=$ $\left[\begin{array}[]{cc}2\Sigma_{Ni}&t_{eff,LR}\\ t_{eff,LR}&2\Sigma_{Bo}\\ \end{array}\right].$ (14)
In this case, the self-energies satisfy $\Sigma_{A(B)}=2\Sigma_{Ni(Bo)}$, which significantly breaks the chiral symmetry of the IF states. For the same-topology configuration, A sites in graphene couple to the same BN species (nitride atoms) on both sides, while B sites couple to boron atoms on both sides. The resulting asymmetric on-site renormalization explains the polarization of the IF charge densities observed in Fig. 4. The eigenstates of $H_{eff,IF,I}$ and $H_{eff,IF,S}$ therefore provide a consistent interpretation of the numerical charge-density distributions.
Figure 4: Charge densities of the interface state $\Sigma_{IF,C}$ in $9_{10}-7_{8}-9_{10}$-AGNRH embedded in BN at $\Delta=2.7$ eV. (a)-(c) Inverse-topology: $t_{in}=0$ ($\Sigma_{IF,C}=7.53$ meV), $t_{in}=0.2~t$ ($\Sigma_{IF,C}=8.81$ meV), $t_{in}=0.8~t$ ($\Sigma_{IF,C}=50.96$ meV). (d)-(f) Same-topology: $t_{in}=0.2~t$ ($\Sigma_{IF,C}=31.35$ meV), $t_{in}=0.4~t$ ($\Sigma_{IF,C}=117.57$ meV), $t_{in}=0.8~t$ ($\Sigma_{IF,C}=0.402$ eV).
[…]
The transmission spectra in the vicinity of the IF resonances are well described by ${\cal T}_{TDQD}(\varepsilon)=\frac{4\Gamma_{e,L}t^{2}_{eff,LR}\Gamma_{e,R}}{|(\varepsilon-E_{L}+i\Gamma_{e,L})(\varepsilon-E_{R}+i\Gamma_{e,R})-t^{2}_{eff,LR}|^{2}}$ in the absence of inter-dot and intra-dot Coulomb interactions [DavidKuo2, 38]. Here, $\Gamma_{e,L(R)}$ denotes the effective tunneling rate between the electrodes and the left (right) TQD, $E_{L(R)}=0$ represent the on-site energies of the two IF states, and $t_{eff,LR}$ is the effective interdot hopping strength. The zero-temperature electrical conductance is given by $G_{e}(\mu)=(2e^{2}/h){\cal T}_{TDQD}(\varepsilon=\mu)$, which reaches the conductance quantum $G_{0}=2e^{2}/h$ when $\mu$ aligns with the energy levels $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$. Here, $\mu$ denotes the Fermi energy of the electrodes.
Notably, the spectra of $G_{e}(\mu)$ indicate that the effective hopping $t_{eff,LR}$ between the IFs in inverse-topology BN environments is enhanced compared with the vacuum boundary condition. This enhancement underscores the constructive role of topology-engineered environments in stabilizing the IFs and strengthening coherent topological transport. In Appendix A, we further investigate the electronic properties of a 15-13-15-AGNRH segment embedded in BN sheets, demonstrating that the above conclusions are not restricted to the 9-7-9 geometry.
[…]
IV Conclusion
In this work, we theoretically investigated the electronic and transport properties of 9-7-9 and 15-13-15 AGNRHs laterally embedded in boron nitride (BN) sheets. Two sandwich configurations were considered: $n$-BNNR/AGNRH/$n$-BNNR and $n$-BNNR/AGNRH/$n$-NBNR, corresponding to same-topology and reverse-topology environments, respectively. These distinct topological configurations give rise to fundamentally different behaviors of the ESs and IFs in the AGNRH.
Within a bulk-boundary perturbation framework, we demonstrate that the topology of the surrounding BN nanoribbons plays a decisive role in determining the stability of the IFs. For same-topology BN environments, interfacial coupling between BN and C atoms induces asymmetric self-energies on the two sublattices. This effect strongly lifts the energy level of the IF state with A-chirality relative to that with B-chirality, breaking the chiral symmetry of the IF subspace and destroying their functionality as TDQDs. As a consequence, the IF levels hybridize with bulk states and lose their spectral isolation.
In contrast, when the surrounding BN nanoribbons realize a reverse-topology configuration, the global inversion symmetry is restored through mirror-inverted interfaces. In this case, the self-energy corrections from the BN environment largely cancel, leaving the IF energy levels only weakly perturbed. The IFs in $n$-BNNR/AGNRH/$n$-NBNR therefore remain well separated from the bulk continuum and retain their TDQD functionality, as confirmed by robust resonant transport reaching the conductance quantum.
Our results establish topology-engineered BN environments as an effective route for stabilizing and even enhancing topological interface states in AGNRHs. The demonstrated symmetry-preserved IFs provide a promising platform for realizing robust quantum functionalities in graphene-based nanoelectronic devices, including high-temperature quantum-dot operation[ChouSY, 39,ChouSY1, 40] and qubit-like applications[LossD, 41Davidkuo5, 43].
Data availability
The data presented in this study are available upon reasonable request.
Conflicts of interest
There are no conflicts to declare
Appendix A Wider width AGNRHs embedded into boron nitride sheet
In the main text, the electronic properties of two sandwich structures, $n$-BNNR/9-7-9-AGNRH/$n$-BNNR and $n$-BNNR/9-7-9-AGNRH/$n$-NBNR, were calculated and analyzed. The former corresponds to a same-topology configuration in which the lower and upper BNNRs share identical topology, whereas the latter represents a reverse-topology configuration. Although only 9-7-9-AGNRH segments have been experimentally synthesized [DJRizzo, 10], in this appendix we investigate the environmental effects on the transport properties of 15-13-15-AGNRHs in order to examine the size dependence of AGNRH responses to BN embedding.
Figure A.1 presents the calculated energy levels of 4-BNNR/15-13-15-AGNRH/4-BNNR and 4-BNNR/15-13-15-AGNRH/4-NBNR segments as functions of the interfacial hopping strength $t_{in}$. The overall trends for the 15-13-15-AGNRH in Fig. A.1 closely resemble those of the 9-7-9 AGNRH shown in Figs. 3(b) and 3(d). However, the bulk band gap of 15-13-15-AGNRH is smaller than that of 9-7-9-AGNRH in both the same-topology and reverse-topology configurations.
In 15-13-15-AGNRHs, six in-gap energy levels appear, consisting of four ESs and two IFs. This follows from the fact that each 15-AGNR segment hosts two left ESs with A-sublattice chirality and two right ESs with B-sublattice chirality. The number of ESs for an $N$-AGNR segment with $R_{1}$ unit cells is given by $N_{es}=(N-3)/6$ and $N_{es}=(N-1)/6$ for $3p$-width and $(3p+1)$-width AGNRs, respectively. For AGNR segments with $R_{2}$ unit cells, the number of ESs becomes $N_{es}+1$. Consequently, the central 13-AGNR segment possesses three ESs at each terminus. The number of IFs at each junction is determined by $N_{IF}=|N_{es,cen}-N_{es,out}|$, where $N_{es,cen}$ and $N_{es,out}$ denote the numbers of ESs in the central and outer AGNR segments, respectively [DavidKuo, 12]. At the 15-13 junction, one IF state with A-chirality emerges, whereas at the 13-15 junction one IF state with B-chirality appears.
In the same topology configuration shown in Fig. A. 1(a), the energy-level separation $\Delta_{IF}$ between $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ for $t_{in}=0$ originates from the wave-function overlap between the A-chirality and B-chirality IFs . We note that $\Delta_{IF}$ in 15-13-15-AGNRH segments is substantially larger than that in 9-7-9-AGNRH segments, reflecting stronger intrinsic coupling between the two IFs. When $t_{in}\leq 0.1~t$, $\Delta_{IF}$ remains nearly unchanged, whereas for $t_{in}>0.2~t$, $\Delta_{IF}$ exhibits an approximate $t_{in}^{2}$ dependence. This behavior is attributed to the gradual decoupling between the A-chirality and B-chirality IFs as their spatial decay lengths are modified (see Fig. A.2(c)).
In the reverse-topology configuration shown in Fig. A.1(b), $\Delta_{IF}=2t_{eff,LR}$ remains nearly constant for $t_{in}\leq 0.3~t$. At $t_{in}=t$, we obtain $E_{g}=0.9$ eV and $\Delta_{IF}=0.362$ eV for the $15_{6}$-$13_{8}$-$15_{6}$-AGNRH segment, compared with $E_{g}=0.978$ eV and $\Delta_{IF}=0.264$ eV for the $9_{6}$-$7_{8}$-$9_{6}$-AGNRH segment. According to Eq. (4), the variation of $\Delta_{IF}$ with respect to $t_{in}$ is governed by the effective inter-IF coupling $t_{eff,LR}$. The results of Fig. A.1(b) therefore indicate that the coherent length of IFs in 15-13-15-AGNRH segments is longer than that in 9-7-9-AGNRH segments, even in the presence of BN embedding.
[…]
In Fig. A.1(a), the energy levels of the topological states are shifted due to the polarization induced by the BN sheets. To further elucidate the polarization effects arising from BNNRs in the same-topology scenario, we plot the charge densities of $\Sigma_{IF,C}$ for the $15_{6}$-$13_{8}$-$15_{6}$ AGNRH at different interfacial hopping strengths $t_{in}$ in Fig. A.2. Figures A.2(a)-2(f) correspond to $t_{in}=0$, $t_{in}=0.2~t$, $t_{in}=0.4~t$, $t_{in}=0.6~t$, $t_{in}=0.8~t$, and $t_{in}=t$, respectively. For $t_{in}=0$ and $t_{in}=0.2~t$, the charge densities of $\Sigma_{IF,C}$ are mostly confined within the graphene ribbon. However, even at $t_{in}=0.2~t$, the A-sublattice sites exhibit a noticeable enhancement of charge density, indicating the onset of BN-induced polarization. When $t_{in}=0.4~t$, the charge density becomes predominantly localized on the A-sublattice sites, which reflects the significant self-energy contributions $\Sigma_{Ni}$ and $\Sigma_{Bo}$ in Eq. (5).
As $t_{in}$ increases further to 0.6t0.6~t, 0.8t0.8~t, and $t$, the charge density of $\Sigma_{IF,C}$ extends to the B-sublattice sites at the BN edges (corresponding to nitride atom positions). Simultaneously, the B-sublattice sites at rows $i=5$ and $i=19$ in the left 15-AGNR segment also display enhanced charge density. This indicates a strong coupling between the A- and B-sublattice sites within the left 15-AGNR segment, resulting in a wave-function delocalization of $\Sigma_{IF,C}$ across the ribbon and into the BN edges.
[…]
To examine the transport properties of the 4-BNNR/15-13-15-AGNRH/4-BNNR segment, we calculate the transmission coefficient ${\cal T}_{GNR}(\varepsilon)$ as a function of applied voltage, which generates a uniform electric field along the negative $y$-direction, as shown in Fig. A.3 for interfacial hopping $t_{in}=0.8~t$ [DingY, 27GSSeal, 29]. The potential induced by the electric field is incorporated into the system Hamiltonian (Eq. (2)) as $U_{y}=-eF_{y}y$, where $F_{y}=V_{y}/L_{a}$, $V_{y}$ is the applied bias, and $L_{a}$ is the length of the AGNRH segment[DavidKuo3, 36,DavidKuo4, 37].
In Fig. A. 3(a) at $V_{y}=0$, two small transmission peaks labeled $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ appear within the bulk gap. Charge transport through these IF levels can be effectively modeled as a single quantum dot (QD) with one energy level and asymmetric tunneling rates $\Gamma_{e,L}\neq\Gamma_{e,R}$ [Davidkuo5, 43]. This demonstrates that the IFs in 4-BNNR/15-13-15-AGNRH/4-BNNR behave as a single TQD at zero bias. Increasing the tunneling rates between the zigzag edge structures of the AGNRH and the electrodes, from $\Gamma_{t}=0.56~\text{eV}$ to $2.7~\text{eV}$, enhances the peak amplitudes, although ${\cal T}_{GNR}$ remains relatively small due to the asymmetry of $\Gamma_{e,L}$ and $\Gamma_{e,R}$, as suggested by the charge density distribution in Fig. A.2(e). As illustrated in Figs. A.3(b)-(d), the separation between the $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ peaks decreases as $V_{y}$ increases. Notably, ${\cal T}_{GNR}$ reaches unity at $V_{y}=1.215$ V with $\Gamma_{t}=0.56~\text{eV}$ in Fig. A.3(d). These results indicate that the 4-BNNR/15-13-15-AGNRH/4-BNNR segment undergoes a transition from a single TQD to a TDQD under electric-field modulation. Finally, we would like to point out that our conclusions remain valid even in the case of $\Delta_{B}\neq-\Delta_{N}$. For example, when the energy level of boron atoms is $\Delta_{B}=2.329~eV$ and the energy level of nitride atoms is $\Delta_{N}=-2.499~eV$, the band inversion symmetry shown in Fig. A.1 is slightly lifted. However, the transmission coefficient still exhibits the resonant tunneling characteristic associated with topological interface states in the 4-BNNR/AGNRH/4-NBNR configuration.
[…]
Acknowledgments
This work was supported by the National Science and Technology Council, Taiwan under Contract No. MOST 107-2112-M-008-023MY2.
E-mail address: mtkuo@ee.ncu.edu.tw